OpenStudy (anonymous):

Given that a truck can drive 25 miles with four gallons of gas, how many gallons does it need to drive 125 miles?

OpenStudy (anonymous):

This is really easy if you think about it :) Ok do you just want the answer or step by step stuff?

OpenStudy (anonymous):

Steps

OpenStudy (anonymous):

what no! @OrangeMaster

OpenStudy (anonymous):

Oh yeah, 20

OpenStudy (poopsiedoodle):

Well, you're taking both variables (gas and miles) and multiplying them by 4. So, it would be 20 gallons.

OpenStudy (anonymous):

125/5=5 5*4=20 gallons.

OpenStudy (anonymous):

Thanks @shoplovepc, I really messed up there.

OpenStudy (anonymous):

Don't I have it right now?

OpenStudy (anonymous):

Ok So first let's look at our FACTS we KNOW 100% that: - the truck can drive 25 miles with four gallons of gas we WANT to know: - how many gallons the truck will need for 125 miles So let's start solving it So we want to know how many times there is 25 in 125, since we know that he needs 4 gallons for 25 miles so 125/25 = 5 And we know that for every 25 miles he needs 4 gallons so 5 * 4 = 20 So he needs 20 gallons for 125 miles.

OpenStudy (anonymous):

Thanks this really helps :)

OpenStudy (anonymous):

yw :)

OpenStudy (poopsiedoodle):

Well, looks like @shoplovepc did a much better job than we did at explaining, @OrangeMaster. Lel

OpenStudy (anonymous):

Thanks again @shoplovepc. I just kept it simple, there wasn't much explaining that needed to be done.

OpenStudy (anonymous):

It was just simple math.

OpenStudy (anonymous):

Any more questions?

OpenStudy (anonymous):

@OrangeMaster haha thanks :) it is pretty simple but it always helps to explain and trains my brain sooooo

OpenStudy (anonymous):

Ya one more question

OpenStudy (anonymous):

At a university, the number of units for a class varies directly as the number of hours each student is expected to work in the class. A four-unit class expects each student to dedicate 10 hours a week to studying for the course. How many hours should a student expect to study for a particular class that is worth six units?

OpenStudy (anonymous):

@OrangeMaster you want to help with this one?

OpenStudy (anonymous):

You have an idea yet, I'm still thinking.

OpenStudy (anonymous):

I'm actually also still thinking

OpenStudy (anonymous):

I think I got it.

OpenStudy (anonymous):

If you divide 10 by 4, you get 2.5

OpenStudy (anonymous):

oh so 15?

OpenStudy (anonymous):

So what I did next is went through numbers until I got 2.5 again. If you divide 15 by 6 you get 2.5 doesn't that make sense.

OpenStudy (anonymous):

Yes, exactly.

OpenStudy (anonymous):

so 15 hours for a six unit class @OrangeMaster smart I can't believe it took me so long to get that

OpenStudy (anonymous):

It took me a minute too, until I figured it out.

OpenStudy (anonymous):

Thanks now I understand

OpenStudy (anonymous):

yw :) also if you have more questions ask ok?

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

lad to help :)

OpenStudy (anonymous):

*Glad